OpenStudy (anonymous):

'Write an equation for the line that passes through (1, 5) and is PERPENDICULAR to the line -7x-5y=18" ?How do I do it?

OpenStudy (amistre64):

same thing ALMOST as a parallel; swap you numbers of x and y; and negate one of them; then fill in the point and change the constant

OpenStudy (amistre64):

-7x-5y=18 ; swap numbers -5x -7y=18 ; negate on of them, ill turn -5 into 5 5x -7y=18 ; now chop off the and and solve with the point given 5(1) -7(5 ) = N

OpenStudy (amistre64):

care to take a try at what N will be?

OpenStudy (anonymous):

-30?

OpenStudy (amistre64):

thats what I get too :) soo.... 5x - 7y = -30 is one way to write the equation of the perp.line

OpenStudy (amistre64):

there are many "forms" that an equation can take, but they all amount to the same line

OpenStudy (anonymous):

I have ALOT of math questions be prepared :)

OpenStudy (amistre64):

for instance: all these are the same value 5x -7y = -30 -5x +7y = 30 y = (5/7)x + 30/7 and a few others